# Inference for Regression with

### 23 Feb 2024 v1

@@ A.2 Proofs for Section 3
$(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\left(T_{3,n}\right)\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\to_{p}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]$
by Assumptions 2(i) and 2(ii).
Now consider $T_{4,n}$. As $\|\hat{\mathbf{p}}_{i}\|\leq 1$, we have
$\|T_{4,n}\|\leq\frac{1}{n}\sum_{i=1}^{n}|\hat{\varepsilon}_{i}^{2}-\varepsilon_{i}^{2}|.$
But note that $\hat{\varepsilon}_{i}-\varepsilon_{i}=\hat{\bm{\theta}}_{i}^{T}(\hat{\bm{\gamma}}-\bm{\gamma})+(\hat{\bm{\theta}}_{i}-\bm{\theta}_{i})^{T}\bm{\gamma}$, where
$\max_{1\leq i\leq n}\left|\hat{\bm{\theta}}_{i}^{T}(\hat{\bm{\gamma}}-\bm{\gamma})\right|\leq\left(\max_{1\leq i\leq n}\|\hat{\bm{\theta}}_{i}-(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\hat{\mathbf{p}}_{i}\|+\|(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\|\right)\|\hat{\bm{\gamma}}-\bm{\gamma}\|\to_{p}0$
@@
by Assumption 2(iii), consistency of $\hat{\bm{\gamma}}$, and the fact that $\|(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\|$ is bounded in probability by Assumptions 2(i) and 2(ii). Moreover,
$\max_{1\leq i\leq n}|(\hat{\bm{\theta}}_{i}-\bm{\theta}_{i})^{T}\bm{\gamma}|\leq\left(\left\|(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}-(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\right\|+\left\|(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\right\|\max_{1\leq i\leq n}\left\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\right\|\right)\left\|\bm{\gamma}\right\|,$
where $(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}-(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\to_{p}0$ by Assumptions 2(i) and 2(ii). Further, as $\hat{\mathbf{p}}_{i}|(C_{i},\bm{\theta}_{i})\sim{C_{i}}^{-1}\mbox{Multinomial}(C_{i},\mathbf{p}_{i})$, for all $t>0$ we have
[…]
by the union bound and Lemma 1 of Mardia et al., (2019). Hence by Assumption 2(v), we have
$\Pr\left(\max_{1\leq i\leq n}\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\|_{1}>t\right)\leq n(2^{V}-2)e^{-\frac{c(\log n)^{1+\epsilon}t^{2}}{2K}},$
for $c,\epsilon>0$. Hence, $\max_{1\leq i\leq n}\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\|_{1}\to_{p}0$ and so $\max_{1\leq i\leq n}|\hat{\varepsilon}_{i}-\varepsilon_{i}|\to_{p}0$. It follows that $\frac{1}{n}\sum_{i=1}^{n}|\hat{\varepsilon}_{i}^{2}-\varepsilon_{i}^{2}|\to_{p}0$ by Hölder’s inequality. Hence, by Assumptions 2(i) and 2(ii) we conclude that $(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\left(T_{4,n}\right)\hat{\mathbf{B}}^{T}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\to_{p}\mathbf{0}$. ∎

### 24 Mar 2024 v2

@@ A.2 Proofs for Section 3
$T_{4,n}=\left[\begin{array}[]{cc}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}(\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}-\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T})&\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}(\hat{\bm{\theta}}_{i}-\bm{\theta}_{i})\mathbf{q}_{i}^{T}\\ \frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\mathbf{q}_{i}(\hat{\bm{\theta}}_{i}-\bm{\theta}_{i})^{T}&\mathbf{0}\end{array}\right].$
Consider the upper-left block. We may deduce by arguments similar to those in the proof of Lemma 6 that
$\left\|\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}(\hat{\bm{\theta}}_{i}\!\!\!\;\hat{\bm{\theta}}_{i}^{T}-\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T})-(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\left(\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}(\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}-\mathbf{p}_{i}\!\!\!\;\mathbf{p}_{i}^{T})\right)\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}\right\|\to_{p}0,$
by Assumption 2(i)-(iii),(v). Since $\mathbf{p}_{i}$ and $\hat{\mathbf{p}}_{i}$ both take values in the simplex, we have $\|\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}-\mathbf{p}_{i}\!\!\!\;\mathbf{p}_{i}^{T}\|\leq 2\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\|$ and so
$\left\|\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}(\hat{\mathbf{p}}_{i}\!\!\!\;\hat{\mathbf{p}}_{i}^{T}-\mathbf{p}_{i}\!\!\!\;\mathbf{p}_{i}^{T})\right\|\leq 2\left(\max_{1\leq i\leq n}\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\|\right)\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\to_{p}0,$
by Lemma 8 and Assumption 2(v). Now consider the lower-left (equivalently, upper-right) block. Again by arguments similar to those in the proof of Lemma 7, we have
$\left\|\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\mathbf{q}_{i}(\hat{\bm{\theta}}_{i}-\bm{\theta}_{i})^{T}-\left(\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\mathbf{q}_{i}(\hat{\mathbf{p}}_{i}-\mathbf{p}_{i})^{T}\right)\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}\right\|\to_{p}0,$
by Assumption 2(i)-(v). But note that
$\left\|\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\mathbf{q}_{i}(\hat{\mathbf{p}}_{i}-\mathbf{p}_{i})^{T}\right\|\leq\left(\max_{1\leq i\leq n}\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\|\right)\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\|\mathbf{q}_{i}\|\to_{p}0,$
by Lemma 8 and Assumption 2(iv)-(v). Therefore, $T_{4,n}\to_{p}\mathbf{0}$.
Now consider $T_{5,n}$. We have
$\hat{\varepsilon}_{i}-\varepsilon_{i}=(\bm{S}\hat{\bm{\theta}}_{i})^{T}(\bm{\gamma}-\hat{\bm{\gamma}})+(\bm{S}(\bm{\theta}_{i}-\hat{\bm{\theta}}_{i}))^{T}\bm{\gamma}+\mathbf{q}_{i}^{T}(\bm{\alpha}-\hat{\bm{\alpha}}),$
where
$\max_{1\leq i\leq n}\left|(\bm{S}\hat{\bm{\theta}}_{i})^{T}(\hat{\bm{\gamma}}-\bm{\gamma})\right|\leq\left(\max_{1\leq i\leq n}\|\bm{S}(\hat{\bm{\theta}}_{i}-(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\hat{\mathbf{p}}_{i})\|+\|\bm{S}(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\|\right)\|\hat{\bm{\gamma}}-\bm{\gamma}\|\to_{p}0$
@@
by Assumption 2(iii), consistency of $\hat{\bm{\gamma}}$, and the fact that $\|(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\|$ is bounded in probability by Assumption 2(i)-(ii) and $\|\hat{\mathbf{p}}_{i}\|\leq 1$. Moreover,
$\max_{1\leq i\leq n}|(\bm{S}(\hat{\bm{\theta}}_{i}-\bm{\theta}_{i}))^{T}\bm{\gamma}|\leq\|\bm{S}\|\bigg{(}\max_{1\leq i\leq n}\|\hat{\bm{\theta}}_{i}-(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}\hat{\mathbf{p}}_{i}\|\\ +\left\|(\hat{\mathbf{B}}\hat{\mathbf{B}}^{T})^{-1}\hat{\mathbf{B}}-(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\right\|+\left\|(\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\right\|\max_{1\leq i\leq n}\left\|\hat{\mathbf{p}}_{i}-\mathbf{p}_{i}\right\|\bigg{)}\left\|\bm{\gamma}\right\|.$
Consider the three terms in parentheses on the right-hand side of this display. The first two terms converge in probability to zero by Assumption 2(i)-(iii) and the third converges in probability to zero by Lemma 8. Finally, we have
$\max_{1\leq i\leq n}\left\|\mathbf{q}_{i}^{T}(\bm{\alpha}-\hat{\bm{\alpha}})\right\|\leq\left(\max_{1\leq i\leq n}\|\mathbf{q}_{i}\|\right)\|\bm{\alpha}-\hat{\bm{\alpha}}\|\to_{p}0$
by $\sqrt{n}$-consistency of $\hat{\bm{\alpha}}$ and the fact that $n^{-1/4}\max_{1\leq i\leq n}\|\mathbf{q}_{i}\|\to_{p}0$ by Assumption 2(iv). Therefore, $\max_{1\leq i\leq n}|\hat{\varepsilon}_{i}-\varepsilon_{i}|\to_{p}0$.
Now, since
$\hat{\varepsilon}_{i}^{2}-\varepsilon_{i}^{2}=2(\hat{\varepsilon}_{i}-\varepsilon_{i})\varepsilon_{i}+(\hat{\varepsilon}_{i}-\varepsilon_{i})^{2},$
we have
$T_{5,n}=\frac{2}{n}\sum_{i=1}^{n}(\hat{\varepsilon}_{i}-\varepsilon_{i})\varepsilon_{i}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}+\frac{1}{n}\sum_{i=1}^{n}(\hat{\varepsilon}_{i}-\varepsilon_{i})^{2}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T},$
and so
$\|T_{5,n}\|$ $\leq 2\left(\max_{1\leq i\leq n}|\hat{\varepsilon}_{i}-\varepsilon_{i}|\right)\frac{1}{n}\sum_{i=1}^{n}|\varepsilon_{i}|\|\hat{\bm{\xi}}_{i}\|^{2}+\left(\max_{1\leq i\leq n}|\hat{\varepsilon}_{i}-\varepsilon_{i}|^{2}\right)\frac{1}{n}\sum_{i=1}^{n}\|\hat{\bm{\xi}}_{i}\|^{2}$
$=\left(\max_{1\leq i\leq n}|\hat{\varepsilon}_{i}-\varepsilon_{i}|\right)\mathrm{tr}\left\{\frac{2}{n}\sum_{i=1}^{n}|\varepsilon_{i}|\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\right\}+\left(\max_{1\leq i\leq n}|\hat{\varepsilon}_{i}-\varepsilon_{i}|^{2}\right)\mathrm{tr}\left\{\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}\right\}\to_{p}0,$
because $\frac{1}{n}\sum_{i=1}^{n}|\varepsilon_{i}|\hat{\bm{\xi}}_{i}\!\!\!\!\;\hat{\bm{\xi}}_{i}^{T}$ is bounded in probability by control of $T_{3,n}$ and $T_{4,n}$, which together imply $\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}^{2}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}$ is bounded in probability, and $\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\xi}}_{i}\!\!\!\;\hat{\bm{\xi}}_{i}^{T}$ is bounded in probability by Lemma 5. ∎
Proof of Theorem 3.
Let $S_{i}=\mathbf{p}_{1,i}\cdot\mathbf{p}_{2,i}$ and $\hat{S}_{i}=\hat{\mathbf{p}}_{1,i}\cdot\hat{\mathbf{p}}_{2,i}$. By standard OLS algebra,
$\sqrt{n}(\hat{\gamma}_{1}-\gamma_{1})=\frac{1}{\frac{1}{n}\sum_{i=1}^{n}(\hat{S}_{i}-\bar{\hat{S}})^{2}}\left(\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\varepsilon_{i}(\hat{S}_{i}-\bar{\hat{S}})-\left(\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(\hat{S}_{i}-S_{i})(\hat{S}_{i}-\bar{\hat{S}})\right)\gamma_{1}\right),$
where $\bar{\hat{S}}=\frac{1}{n}\sum_{i=1}^{n}\hat{S}_{i}$.
By Chebyshev’s inequality, for all integers $k_{1},k_{2}\geq 0$ and all $t>0$, we have
$\Pr\left(\left|\frac{1}{n}\sum_{i=1}^{n}\hat{S}_{i}^{k_{1}}S_{i}^{k_{2}}-\mathbb{E}\left[\hat{S}_{i}^{k_{1}}S_{i}^{k_{2}}\right]\right|>t\right)\leq\frac{1}{nt^{2}}\mathbb{E}\left[\hat{S}_{i}^{2k_{1}}S_{i}^{2k_{2}}\right]\leq\frac{1}{nt^{2}},$ (26)
because $|\hat{S}_{i}|\leq\|\hat{\mathbf{p}}_{1,i}\|\|\hat{\mathbf{p}}_{2,i}\|\leq 1$ by virtue of the fact that $\|\hat{\mathbf{p}}_{t,i}\|\leq\|\hat{\mathbf{p}}_{t,i}\|_{1}=1$ for $t=1,2$, with $\|\cdot\|_{1}$ denoting the $\ell^{1}$ norm, and similarly for $S_{i}$. Let $F_{i}=(\mathbf{p}_{1,i},\mathbf{p}_{2,i},C_{1,i},C_{2,i})$. Note that
$\mathbb{E}\left[\left.\hat{S}_{i}\right|F_{i}\right]=S_{i}$
because $\mathbf{x}_{1,i}$, and $\mathbf{x}_{2,i}$ are independent conditional on $(C_{1,i},C_{2,i},\mathbf{p}_{1,i},\mathbf{p}_{2,i})$. Hence, $\mathbb{E}[\hat{S}_{i}]=\mathbb{E}[S_{i}]$ and so it follows by (26) that $\bar{\hat{S}}\to_{p}\mathbb{E}[S_{i}]$. Moreover, for any conformable non-stochastic matrix $\mathbf{M}$, we have for $t=1,2$ that
$\mathbb{E}\left[\left.\hat{\mathbf{p}}_{t,i}^{T}\mathbf{M}\hat{\mathbf{p}}_{t,i}\right|F_{i}\right]=\mathbf{p}_{t,i}^{T}\mathbf{M}\mathbf{p}_{t,i}+\frac{1}{C_{t,i}}\,\mathrm{tr}\left\{\mathbf{M}\left(\mathrm{diag}(\mathbf{p}_{t,i})-\mathbf{p}_{t,i}\mathbf{p}_{t,i}^{T}\right)\right\}.$
Hence, by independence of $\mathbf{x}_{1,i}$ and $\mathbf{x}_{2,i}$ conditional on $F_{i}$, we have
$\mathbb{E}\left[\left.\hat{S}_{i}^{2}\right|F_{i}\right]$ $=\mathbb{E}\left[\left.\mathbb{E}\left[\left.\hat{\mathbf{p}}_{1,i}^{T}(\hat{\mathbf{p}}_{2,i}\hat{\mathbf{p}}_{2,i}^{T})\hat{\mathbf{p}}_{1,i}\right|\hat{\mathbf{p}}_{2,i},F_{i}\right]\right|F_{i}\right]$
$=\mathbb{E}\left[\left.\mathbf{p}_{1,i}^{T}(\hat{\mathbf{p}}_{2,i}\hat{\mathbf{p}}_{2,i}^{T})\mathbf{p}_{1,i}+\frac{1}{C_{1,i}}\,\mathrm{tr}\left\{(\hat{\mathbf{p}}_{2,i}\hat{\mathbf{p}}_{2,i}^{T})\left(\mathrm{diag}(\mathbf{p}_{1,i})-\mathbf{p}_{1,i}\mathbf{p}_{1,i}^{T}\right)\right\}\right|F_{i}\right]$
$=\mathbb{E}\left[\left.\hat{\mathbf{p}}_{2,i}^{T}(\mathbf{p}_{1,i}\mathbf{p}_{1,i}^{T})\hat{\mathbf{p}}_{2,i}\right|F_{i}\right]+\frac{1}{C_{1,i}}\mathbb{E}\left[\left.\hat{\mathbf{p}}_{2,i}^{T}\left(\mathrm{diag}(\mathbf{p}_{1,i})-\mathbf{p}_{1,i}\mathbf{p}_{1,i}^{T}\right)\hat{\mathbf{p}}_{2,i}\right|F_{i}\right]$
$=S_{i}^{2}+\frac{1}{C_{2,i}}\mathbf{p}_{1,i}^{T}\left(\mathrm{diag}(\mathbf{p}_{2,i})-\mathbf{p}_{2,i}\mathbf{p}_{2,i}^{T}\right)\mathbf{p}_{1,i}+\frac{1}{C_{1,i}}\mathbf{p}_{2,i}^{T}\left(\mathrm{diag}(\mathbf{p}_{1,i})-\mathbf{p}_{1,i}\mathbf{p}_{1,i}^{T}\right)\mathbf{p}_{2,i}$
$\quad\quad+\frac{1}{C_{1,i}C_{2,i}}\mathrm{tr}\left\{\left(\mathrm{diag}(\mathbf{p}_{2,i})-\mathbf{p}_{2,i}\mathbf{p}_{2,i}^{T}\right)\left(\mathrm{diag}(\mathbf{p}_{1,i})-\mathbf{p}_{1,i}\mathbf{p}_{1,i}^{T}\right)\right\}.$ (27)
It follows by (11) and (26) that
$\left|\frac{1}{n}\sum_{i=1}^{n}\hat{S}_{i}^{2}-\mathbb{E}[S_{i}^{2}]\right|\to_{p}0.$
Hence, $\frac{1}{n}\sum_{i=1}^{n}(\hat{S}_{i}-\bar{\hat{S}})^{2}\to_{p}\mathrm{Var}(S_{i})$.
For the numerator term, note that
$\left|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\varepsilon_{i}(\hat{S}_{i}-\bar{\hat{S}})-\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\varepsilon_{i}(S_{i}-\mathbb{E}[S_{i}])\right|\to_{p}0,$
because: firstly,
$\Pr\left(\left|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\varepsilon_{i}(\hat{S}_{i}-S_{i})\right|>t\right)\leq\frac{1}{t^{2}}\mathbb{E}\left[\mathbb{E}\left[\left.\varepsilon_{i}^{2}\right|F_{i}\right]\mathbb{E}\left[\left.\hat{S}_{i}^{2}-S_{i}^{2}\right|F_{i}\right]\right]\to 0$
by (11) and (27), mutual independence of $C_{1,i}$, $C_{2,i}$, and $(\mathbf{p}_{1,i},\mathbf{p}_{2,i},Y_{i})$, and using finite second moment of $\varepsilon_{i}$; and, second,
$\left|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\varepsilon_{i}(\bar{\hat{S}}-\mathbb{E}[S_{i}])\right|\leq\left|\bar{\hat{S}}-\mathbb{E}[S_{i}]\right|\left|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\varepsilon_{i}\right|\to_{p}0$
by the CLT and consistency of $\bar{\hat{S}}$. So by a second application of the CLT we deduce that
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\varepsilon_{i}(\hat{S}_{i}-\bar{\hat{S}})\to_{d}N\left(0,\mathbb{E}\left[\varepsilon_{i}^{2}(S_{i}-\mathbb{E}[S_{i}])^{2}\right]\right).$
Finally to characterize the bias term, first note that
$\left|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(\hat{S}_{i}-S_{i})(\hat{S}_{i}-\bar{\hat{S}})-\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(\hat{S}_{i}-S_{i})\hat{S}_{i}\right|=|\bar{\hat{S}}|\left|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(\hat{S}_{i}-S_{i})\right|\to_{p}0$
by consistency of $\bar{\hat{S}}$ and because
$\mathbb{E}[(\hat{S}_{i}-S_{i})^{2}]\to 0$
holds by (11) and (27) and mutual independence of $C_{1,i}$, $C_{2,i}$, and $(\mathbf{p}_{1,i},\mathbf{p}_{2,i},Y_{i})$. Hence by Chebyshev’s inequality, we have
$\Pr\left(\left|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left((\hat{S}_{i}-S_{i})\hat{S}_{i}-\mathbb{E}\left[(\hat{S}_{i}-S_{i})\hat{S}_{i}\right]\right)\right|>t\right)\leq\frac{1}{t^{2}}\mathbb{E}[(\hat{S}_{i}-S_{i})^{2}\hat{S}_{i}^{2}]\\ \leq\frac{1}{t^{2}}\mathbb{E}[(\hat{S}_{i}-S_{i})^{2}]\to 0,$
where the second inequality is because $|\hat{S}_{i}|\leq 1$. We have therefore shown that
$\left|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(\hat{S}_{i}-S_{i})(\hat{S}_{i}-\bar{\hat{S}})-\mathbb{E}[\sqrt{n}(\hat{S}_{i}-S_{i})\hat{S}_{i}]\right|\to_{p}0.$
Finally,
$\mathbb{E}[\sqrt{n}(\hat{S}_{i}-S_{i})\hat{S}_{i}]\to\kappa_{1}\mathbb{E}\left[\mathbf{p}_{2,i}^{T}\left(\mathrm{diag}(\mathbf{p}_{1,i})-\mathbf{p}_{1,i}\mathbf{p}_{1,i}^{T}\right)\mathbf{p}_{2,i}\right]\\ +\kappa_{2}\mathbb{E}\left[\mathbf{p}_{1,i}^{T}\left(\mathrm{diag}(\mathbf{p}_{2,i})-\mathbf{p}_{2,i}\mathbf{p}_{2,i}^{T}\right)\mathbf{p}_{1,i}\right],$
by (11), (27), and mutual independence of $C_{1,i}$, $C_{2,i}$, and $(\mathbf{p}_{1,i},\mathbf{p}_{2,i},Y_{i})$. ∎